# QGait: Toward Accurate Quantization for Gait Recognition with Binarized Input

### v1

@@ 3.3 Inter-class Distance-guided Distillation
For the well-refined low-bit model, we assume that the KL divergence loss is minimized, aligning the predicted probability density with that of the high-bit model. For any pair of class indices , it gives:
@@
Inspired by the remarkable performance of triplet loss in gait recognition [29; 11], we visualize the distribution of embedding vectors from different samples as shown in Fig. 4. Both low-bit and high-bit networks can effectively cluster samples of the same class. Despite minor differences in inner-class distributions, the impact of quantization-induced numerical disparities on performance is negligible compared to noticeable changes in inter-class distances (dashed lines in Fig. 4). Based on the observation of heterogeneous distances, we decide to approach distillation from the perspective of distance distribution. Considering a mini-batch of data, the probability vectors guided by the self-similarity of distances between samples’ embeddings are defined as:
@@ 4 Experimental Results
To validate the generalization and effectiveness of our approach, we performed comprehensive experiments on the mainstream datasets and compared our approach with state-of-the-art gait methods and quantization methods. We selected representative methods Dorefa[45], PACT[6], and LSQ[9] as the baselines for quantization. QGait models in subsection 4.2 are implemented based on GaitBase[11]. Other quantized models are reported in subsection 4.4. The quantization bit-width is set to 4-bit or 8-bit, as in practical scenarios, devices often do not support other types of bits. The code implementation is based on PyTorch 2.0, and all experiments were conducted using RTX A6000s.
GaitSet [4] | 32 | 10.68 | 36.75 | 58.66 | 64.20 | 30.21 | 17.55 |
GaitPart [12] | 32 | 10.67 | 28.44 | 47.58 | 53.70 | 22.01 | 12.32 |
GaitGL [22] | 32 | 41.64 | 29.88 | 48.69 | 54.42 | 22.30 | 13.50 |
DyGait [37] | 32 | 652.22 | 66.30 | 80.80 | 86.10 | 56.40 | 37.30 |
GaitBase [11] | 32 | 118.30 | 64.40 | 81.50 | 85.80 | 55.28 | 36.73 |
Dorefa-4 [45] | w4/a4 | - | 60.00 | 78.30 | 83.20 | 50.32 | 32.66 |
PACT-4 [6] | w4/a4 | - | 59.80 | 78.00 | 82.80 | 49.56 | 31.53 |
LSQ-4 [9] | w4/a4 | - | 63.80 | 79.60 | 86.30 | 54.22 | 35.71 |
QGait-4 | w4/a4 | 1.85 | 64.60 | 81.20 | 86.20 | 54.29 | 35.76 |
Dorefa-8 [45] | w8/a8 | - | 64.30 | 80.80 | 86.10 | 54.34 | 35.87 |
PACT-8 [6] | w8/a8 | - | 63.30 | 80.50 | 85.20 | 53.41 | 34.64 |
LSQ-8 [9] | w8/a8 | - | 64.80 | 81.40 | 85.90 | 55.15 | 36.76 |
QGait-8 | w8/a8 | 7.39 | 66.50 | 81.30 | 86.20 | 55.82 | 37.05 |
@@ 4.1 Experimental Settings
We pre-process the original gait sequences for all the datasets following previous works [11; 37; 10] and the size of each frame is set to 64 × 44. For Gait3D[43], GREW[46] and OUMVLP[33], the number of Identities and the number of samples per identity is set to and for CASIA-B [40]. The optimizer is Adam with a learning rate of 1e-4 and momentum of 0.9 for all the datasets for a fair comparison. The frame number settings are consistent with OpenGait[11]. In the first training stage of QGait, the iterations are set to 60K, 180K, and 120Krespectively for Gait3D[43], GREW[46] and OUMVLP[33]. In the fine-tuning stage, we only conduct 1K or 2K iterations according to different datasets to avoid overfitting while ensuring a fair comparison relative to the full-precision model.
@@
We evaluated on the test set corresponding to the training dataset. In addition to comparing with quantized models, we also included comparisons with state-of-the-art full-precision models, including GaitSet[4], GaitPart[12], GaitGL[22], DyGait[36] and GaitBase[11]. Evaluation metrics used include Rank-n, mAP (mean Average Precision), mINP (mean Inverted Normalized Precision), and BitOPs (Bit Operations Per Second). Conventionally, given the weight of and input feature of , BitOPs of a quantized convolution layer can be calculated as . The BitOPs of each model on different datasets are calculated based on different frames (100 frames for Gait3D and OUMVLP, 120 frames for GREW), along with different model configurations.
@@
GaitSet [4] | 32 | 50.68 | 46.30 | 63.62 |
GaitPart [12] | 32 | 31.68 | 43.99 | 60.74 |
GaitGL [22] | 32 | 234.21 | 47.28 | 63.88 |
DyGait [37] | 32 | 867.14 | 71.40 | 83.21 |
GaitBase [11] | 32 | 141.96 | 60.12 | 75.47 |
Dorefa-4 [45] | w4/a4 | - | 53.31 | 69.68 |
PACT-4 [6] | w4/a4 | - | 52.40 | 68.72 |
LSQ-4 [9] | w4/a4 | - | 57.44 | 73.51 |
QGait-4 | w4/a4 | 2.22 | 58.52 | 74.06 |
Dorefa-8 [45] | w8/a8 | - | 60.00 | 74.93 |
PACT-8 [6] | w8/a8 | - | 59.03 | 73.62 |
LSQ-8 [9] | w8/a8 | - | 59.89 | 75.13 |
QGait-8 | w8/a8 | 8.87 | 60.60 | 75.89 |
Methods | Bit | BitOPs | Rank-1 | Rank-5 |
GaitSet [4] | 32 | 42.29 | 87.12 | 91.64 |
GaitPart [12] | 32 | 26.44 | 88.68 | 92.07 |
GaitGL [22] | 32 | 193.30 | 89.73 | 92.32 |
DyGait [37] | 32 | 829.79 | 90.47 | 92.89 |
GaitBase [11] | 32 | 118.30 | 90.32 | 92.67 |
Dorefa-4 [45] | w4/a4 | - | 88.24 | 92.02 |
PACT-4 [6] | w4/a4 | - | 87.65 | 91.96 |
LSQ-4 [9] | w4/a4 | - | 86.19 | 91.19 |
@@ 4.2 Comparison with State-of-the-art Methods
Evaluation on Gait3D
The Gait3D dataset is a large-scale dataset, which contains 4,000 subjects and over 25,000 sequences captured from an unconstrained indoor scene by 39 cameras. The diversity and complexity of the data make many models perform poorly on this dataset. As Tab. 1 shows, in the 32-bit full-precision model, DyGait achieves state-of-the-art performance, albeit at the expense of costly BitOPs. While GaitSet and GaitPart have computational advantages, their accuracy falls short of requirements. QGait achieves performance over the full-precision model with very low BitOPs and achieves precision lossless compared to other quantization methods. The QGait-8 outperforms GaitBase with 2.1% increase of Rank-1 accuracy. QGait-4 surpasses other quantization methods and even outperforms the full-precision model in Rank-1 accuracy. For complex datasets like Gait3D, the introduction of quantization may act as a form of regularization, enhancing the network’s generalization ability, thus achieving significant performance improvement.
Evaluation on GREW
The GREW dataset is a large-scale outdoor gait dataset. It can be observed from Tab. 3(a) that GaitGL, despite its high computational costs (with BitOPs exceeding 200G), does not achieve the expected high performance. For 8-bit quantized models, only QGait achieves lossless compression relative to the full-precision model. While QGait surpasses other quantization methods at 4-bit, e.g., QGait achieves a Rank-1 accuracy exceeding LSQ by 1.08%, it still incurs a loss in accuracy. This may be because exploration of large-scale outdoor datasets like GREW is still insufficient in the full-precision model, thus the errors introduced by quantization cannot be effectively compensated for through training.
Evaluation on OUMVLP
Although the OUMVLP dataset contains a large amount of data, the gait sequences in this dataset are all in normal states without occlusions, disguises, or other variations. From Tab. 3(b), it can be observed that QGait incurs some loss in accuracy at 4-bit, while there is minimal loss at 8-bit. This suggests that using stronger baseline models may lead to better performance on the OUMVLP dataset. Nevertheless, compared to other quantization methods, we still maintain a lead of over 1% in accuracy at 4-bit.
@@ 4.3 Ablation Study
Baseline | Vanilla KD | IDD | FT | Rank-1 (%) |
@@
Training with Soft Quantizer We empirically demonstrate that introducing a soft quantizer at the beginning of training indeed significantly affects the model’s convergence. As the parameter k increases, the soft quantizer approaches the round function more closely, resulting in smaller updates to gradients at each step, as illustrated in Fig.5 in Appendix A.2. We also experimentally determine how to update the value during the fine-tuning stage. After trying several different strategies, we decided to use a gradually increasing approach, where we increment the value of k as the number of iterations increases until it reaches a predetermined threshold as Tab. 4(a). The selection of the threshold is based on the results calculated in the methodology section.
@@
Inter-class Distance-guided Distillation To study the effects of IDD together with other components, we conducted module ablation experiments on the baseline method LSQ. It can be observed from Tab. 4(b) that, as discussed in the methodology section, vanilla KD overly balances the quantization error, resulting in decreased accuracy. IDD improves the accuracy by 1.84%, and even after incorporating the soft quantizer fine-tuning strategy, it still achieves a further 0.86% improvement in accuracy.
@@ 4.4 Quantization of Different Models
We also conduct experiments on different model architectures using our proposed QGait method, including GaitSet, GaitPart and DyGait. It can be seen from Tab. 4 that QGait shows comparable performance with full-precision models on different model structures. For two lightweight models, the 4-bit quantized models lost some precision. However, the quantized 4-bit DyGait model outperforms the full-precision model. That is because the quantization acts as a regularization. It can alleviate overfitting on some heavy models while the data is relatively simple. This property may inspire the later works.
@@ 5 Conclusion
In this paper, we explore for the first time a gait recognition network based on quantization compression. Addressing the data characteristics of gait recognition tasks and the difficulty of training convergence, we design quantization schemes that are different from conventional tasks. Experimental results demonstrate that the two proposed methods effectively improve the quantization accuracy of gait recognition. In experiments, we found that different datasets exhibit varying levels of robustness to quantization, which may be related to the complexity of the data. This brings new perspectives and considerations for future research in gait recognition.
Appendix A Appendix
A.1 Calculation of Integral
The methods section of this paper involves the calculation of two key integrals. The following is the derivation of the corresponding two indefinite integrals.
Let be the integral variable, the integral of can be calculated as:
For the integrand , substitute and , we get:
For the integrand , substitute and let , then we have:
where is a constant. Substitute back for and , we get:
A.2 Direct Training with Soft Quantizer
The parameter controls the steepness of the curve. As increases, the curve becomes closer to the curve of the round function. In the body part, We have demonstrated through theoretical analysis that the convergence of is worse than using STE. Here, we use different values as parameters for the soft quantizer, which are actually applied to network training. The results are as Fig. 5 shows. It is obvious that as increases, the convergence speed of the network becomes very slow. Although the soft quantizer simulates quantization errors, the decrease in gradient mean and the increase in variance make learning task losses exceptionally difficult.
A.3 Limitations
While quantization can bring actual acceleration effect to gait recognition networks, there are also some limitations:
Model Deployment: In the training stage, the model simulates the quantization error with pure floating-point numbers so as to obtain the corresponding network weight which can adapt to the error. In the inference phase, it is necessary to select the appropriate strategy for deployment according to the different deployment platforms (such as GPU and mobile). This process may involve model optimization, model transformation, model quantization, and model compilation optimization. Maintaining alignment of model accuracy during this process requires considerable experience.
Quantization of Different Models: The structural differences between different models can be significant. Quantization Aware Training on these structures requires manual modifications to the model files. At the same time, the influence of different topologies on quantization is unknown. The scalability of quantization needs to be based on the understanding of model properties and the adjustment of quantization parameters.
Hardware Support: Quantization requires specific hardware support to achieve acceleration. For example, the NVIDIA Tesla V100 graphics card supports FP32, FP16, INT8 and INT4 precision. The A100 graphics card in addition to supporting FP32 and FP16, also introduces new precision TF32 and FP64. The A100 also supports mixed precision BF16/FP16 as well as INT8, INT4, and Binary.
A.4 Broader Impacts
The quantization and compression of gait recognition networks have the potential to bring about various societal impacts, both positive and negative. Here is a detailed discussion of these impacts:
Increased Portability and Usability: Quantization and compression enable gait recognition models to be deployed on resource-constrained devices such as smartphones and IoT devices, facilitating the use of gait recognition technology in mobile and embedded systems. Reduced model size and computational requirements make real-time processing feasible, enhancing the responsiveness and user experience of gait recognition systems.
Energy Savings and Cost Reduction: Compressed models demand less computational power, thereby reducing energy consumption, which is particularly beneficial for data centers and edge devices, contributing to a greener computing environment. Lower hardware and maintenance costs due to reduced computational demands make gait recognition technology more affordable for businesses and consumers, promoting wider adoption.
Enhanced Privacy Protection: Quantized and compressed models are well-suited for running on local devices, which minimizes the need to transmit data to the cloud. This local processing helps protect sensitive information as data does not leave the device.
However, there are also potential negative impacts:
Privacy and Security Concerns: The misuse of gait recognition technology can lead to privacy infringements. Governments, companies, or individuals could potentially exploit this technology for extensive surveillance and unauthorized monitoring. As the technology becomes more ubiquitous due to easier deployment via quantized models, the risk of unauthorized data collection and analysis increases, raising privacy concerns.
Technical Limitations and Reliability Issues: Despite efforts to maintain accuracy, quantized models might suffer from reduced precision, particularly in complex scenarios or when dealing with diverse populations, which can lead to misidentification or non-recognition. Compressed models may exhibit lower adaptability to new data or varied environments, necessitating frequent updates and optimizations.
Social and Ethical Issues: The widespread application of gait recognition raises ethical questions related to biometric technology, such as transparency, informed consent, and user rights. These issues might be exacerbated with the proliferation of quantized models due to lower barriers to entry. Additionally, if the training datasets are biased, gait recognition systems could inherit these biases, leading to inconsistent recognition rates across different groups, thus causing discrimination issues.

### v2

@@ 3.3 Inter-class Distance-guided Calibration
For the well-trained low-bit model, we assume that the KL divergence loss is minimized, aligning the predicted probability density with that of the high-bit model. For any pair of class indices , it gives:
@@
Inspired by the remarkable performance of triplet loss in gait recognition [Schroff2015FaceNetAU, Fan2022OpenGaitRG], we visualize the distribution of embedding vectors from different samples as shown in Fig. 5. Both low-bit and high-bit networks can effectively cluster samples of the same class. Despite minor differences in inner-class distributions, the impact of quantization-induced numerical disparities on performance is negligible compared to noticeable changes in inter-class distances (dashed lines in Fig. 5). Based on the observation of heterogeneous distances, we decided to approach a calibration from the perspective of distance distribution. Considering a mini-batch of data, the probability vectors guided by the self-similarity of distances between samples’ embeddings are defined as:
@@ 4 Experimental Results
To validate the generalization and effectiveness of our approach, we performed comprehensive experiments on the mainstream datasets and compared our approach with state-of-the-art gait methods and quantization methods. We selected representative methods Dorefa [Zhou2016DoReFaNetTL], PACT [Choi2018PACTPC], and LSQ [Esser2019LearnedSS] as the baselines for quantization. QGait models in subsection 4.2 are implemented based on GaitBase [Fan2022OpenGaitRG]. Other quantized models are reported in subsection 4.4. The code implementation is based on PyTorch 2.0, and all experiments were conducted using RTX A6000s.
| Methods | Bit | BitOPs | Rank-1 | Rank-5 |
| GaitSet | 32 | 50.68 | 46.30 | 63.62 |
| GaitPart | 32 | 31.68 | 43.99 | 60.74 |
| GaitGL | 32 | 234.21 | 47.28 | 63.88 |
| DyGait | 32 | 867.14 | 71.40 | 83.21 |
| GaitBase | 32 | 141.96 | 60.12 | 75.47 |
| Dorefa-4 | w4/a4 | - | 53.31 | 69.68 |
| PACT-4 | w4/a4 | - | 52.40 | 68.72 |
| LSQ-4 | w4/a4 | - | 57.44 | 73.51 |
| QGait-4 | w4/a4 | 2.22 | 58.52 | 74.06 |
| Dorefa-8 | w8/a8 | - | 60.00 | 74.93 |
| PACT-8 | w8/a8 | - | 59.03 | 73.62 |
| LSQ-8 | w8/a8 | - | 59.89 | 75.13 |
| QGait-8 | w8/a8 | 8.87 | 60.60 | 75.89 |
@@ 4.1 Experimental Settings
We pre-process the original gait sequences for all the datasets following previous works [Fan2022OpenGaitRG, Wang2023DyGaitED, Fan2023ExploringDM] and the size of each frame is set to 64 × 44. For Gait3D [Zheng2022GaitRI], GREW [Zhu2021GaitRI] and OUMVLP [Takemura2018MultiviewLP], the number of Identities and the number of samples per identity is set to and for CASIA-B [Yu2006AFF]. The optimizer is Adam with a learning rate of 1e-4 and momentum of 0.9 for all the datasets for a fair comparison. The frame number settings are consistent with OpenGait [Fan2022OpenGaitRG]. In the first training stage of QGait, the iterations are set to 60K, 180K, and 120Krespectively for Gait3D [Zheng2022GaitRI], GREW [Zhu2021GaitRI] and OUMVLP [Takemura2018MultiviewLP]. In the fine-tuning stage, we only conduct 1K or 2K iterations according to different datasets to avoid overfitting while ensuring a fair comparison relative to the full-precision model.
@@
We conducted the evaluation on the test set corresponding to the training dataset. In addition to comparing with quantized models, we also included comparisons with state-of-the-art full-precision models, including GaitSet [Chao2018GaitSetRG], GaitPart [Fan2020GaitPartTP], GaitGL [Lin2020GaitRV], DyGait [Wagg2004OnAM] and GaitBase [Fan2022OpenGaitRG]. Evaluation metrics used include Rank-n, mAP (mean Average Precision), mINP (mean Inverted Normalized Precision), and BitOPs (Bit Operations Per Second). Conventionally, given the weight of -bit and input feature of -bit, BitOPs of a quantized convolution layer can be calculated as . The BitOPs of each model on different datasets are calculated based on different frames (100 frames for Gait3D and OUMVLP, 120 frames for GREW), along with different model configurations.
@@
| GaitSet | 32 | 42.29 | 87.12 | 91.64 |
| GaitPart | 32 | 26.44 | 88.68 | 92.07 |
| GaitGL | 32 | 193.30 | 89.73 | 92.32 |
| DyGait | 32 | 829.79 | 90.47 | 92.89 |
| GaitBase | 32 | 118.30 | 90.32 | 92.67 |
| Dorefa-4 | w4/a4 | - | 88.24 | 92.02 |
| PACT-4 | w4/a4 | - | 87.65 | 91.96 |
| LSQ-4 | w4/a4 | - | 86.19 | 91.19 |
@@ 4.2 Comparison with State-of-the-art Methods
Figure 1, Fig. 2, and Fig. 3 respectively provide comparisons between QGait and various SOTA models on different datasets. QGait achieves the optimal balance between model performance and computational complexity. For complex datasets like Gait3D, the introduction of quantization may act as a form of regularization, enhancing the network’s generalization ability.
@@ 4.3 Ablation Study
| Baseline | Vanilla KD | IDC | FT | Rank-1 (%) |
@@
Training with Soft Quantizer
We empirically demonstrate that introducing a soft quantizer at the beginning of training indeed significantly affects the model’s convergence. As the parameter increases, the soft quantizer approaches the round function more closely, resulting in smaller updates to gradients at each step as Fig. 6 shows. We also experimentally determine how to update the value during the fine-tuning stage. After trying several different strategies, we decided to use a gradually increasing approach, where we increment the value of k as the number of iterations increases until it reaches a predetermined threshold as Fig. 4 shows.
Inter-class Distance-guided Calibration
To study the effects of IDC together with other components, we conducted module ablation experiments on the baseline method LSQ. It can be observed from Fig. 5 that, as discussed in the methodology section, vanilla KD overly balances the quantization error, resulting in decreased accuracy. IDC improves the accuracy by 1.84%, and even after incorporating the soft quantizer fine-tuning strategy, it still achieves a further 0.86% improvement in accuracy.
@@ 4.4 Quantization of Different Models
We also conduct experiments on different model architectures using our proposed QGait method, including GaitSet, GaitPart and DyGait. It can be seen from Fig. 6 that QGait shows comparable performance with full-precision models on different model structures. For two lightweight models, the 4-bit quantized models lost some precision. However, the quantized 4-bit DyGait model outperforms the full-precision model. That is because the quantization acts as a regularization. It can alleviate overfitting on some heavy models while the data is relatively simple. This property may inspire the later works.
4.5 Quantization of Different Backbones
Future gait models may require larger-scale backbones, rather than being limited to just ResNet9. To demonstrate the effectiveness of quantization, we individually quantized different backbones to observe their latency as Fig. 7 shows.
| Bit | Params | BitOPs | Latency | |
| ResNet-18 | 32 | 44.6M | 1858G | 18.2ms |
| ResNet-18 | 8 | 11.1M | 116G | 5.7ms |
| ResNet-18 | 4 | 5.8M | 34G | 4.1ms |
| ResNet-50 | 32 | 97.8M | 6951G | 50.6ms |
| ResNet-50 | 8 | 24.5M | 247G | 15.9ms |
| ResNet-50 | 4 | 13.1M | 67G | 11.4ms |
@@ 5 Conclusion
In this paper, we explore a gait recognition network based on quantization compression for the first time. Addressing the data characteristics of gait recognition tasks and the difficulty of training convergence, we design quantization schemes that are different from conventional tasks. Experimental results demonstrate that the two proposed methods effectively improve the quantization accuracy of gait recognition. In experiments, we found that different datasets exhibit varying levels of robustness to quantization, which may be related to the complexity of the data. This brings new perspectives and considerations for future research in gait recognition.